OpenStudy (anonymous):

The following equation has an unknown (a) as one of its coefficients for x. 7x-9x-10=2(x-4)-ax What value for a ensures that this equation has no solutions ?

OpenStudy (anonymous):

alright so 1st we will simplify this equation \(7x-9x-10=2(x-4)-ax \)

OpenStudy (anonymous):

\(7x-9x-10=2(x-4)-ax \) \(-2x-10=2x-8-ax\) \(-2x-10+2x=2x-8-ax+2x\) \(-10=-8+4x-ax\) now if the 'a' becomes 4 then- \(-10=-8+4x-4x\) \(-10=-8\) here this thing -> -10=-8 is false so whatever value of x you take it will not satisfy the equation so a=4 is correct :)

OpenStudy (anonymous):

@jaylajohnson

OpenStudy (anonymous):

Thank you so much for your help @paperairplanes